# TTS-VAR: A Test-Time Scaling Framework for Visual Auto-Regressive Generation

### v1

@@ D.3 $\lambda$ Setting
Appendix E More Visualization Results
We present pairs of results on GenEval in Fig. 9 to compare the quality of Infinity, Infinity-IS, Infinity-BoN, and Infinity-TTS-VAR. These cases are sampled from text prompts in GenEval, which include the single object, two objects, counting, colors, position, and color attributes. As shown, our method produces higher-quality samples for the single object, such as the airplane in the left second line, effectively avoiding the generation of artifacts. In two-object settings, TTS-VAR successfully distinguishes references to different objects and generates accurate outputs based on prompts. For example, in the right second line, it eliminates conceptual mixtures and object disappearances. As illustrated in lines 3-10 on the right side, TTS-VAR also excels at determining counting numbers, positional relationships, and color attributes. Notably, in the last right line, our method generates a counterintuitive "green carrot," demonstrating its ability to separate objects from their natural attributes.
@@ Appendix E More Visualization Results
Appendix F Societal Impact
@@
Appendix G Limitation and Future Work
Though TTS-VAR shows significant improvement over the baseline and sets a new record, it still has two main limitations. First, TTS-VAR does not completely address the misalignment between text prompts and generated images. As indicated by the scores in Table 7, there are still some failure cases, particularly in the Position item. Second, while TTS-VAR is based on a general coarse-to-fine process, its potential application to other coarse-to-fine models, such as autoregressive models that use 1-D tokenizers, remains unexplored. In the future, we will investigate the generation process more thoroughly, examining the reasons for failure and designing solutions to unlock further scaling potential. Additionally, we plan to assess the compatibility of TTS-VAR with other autoregressive coarse-to-fine models, including those utilizing 1-D tokenizers and hybrid architectures that combine diffusion models. These efforts aim to create a more robust scaling framework for text-to-image synthesis while enhancing the methodological transferability of coarse-to-fine paradigms.

### v2

@@ G.5 $\lambda$ Setting
Appendix H More Visualization Results
We present pairs of results on GenEval in Fig.˜9 to compare the quality of Infinity, Infinity-IS, Infinity-BoN, and Infinity-TTS-VAR. These cases are sampled from text prompts in GenEval, which include the single object, two objects, counting, colors, position, and color attributes. As shown, our method produces higher-quality samples for the single object, such as the airplane in the left second line, effectively avoiding the generation of artifacts. In two-object settings, TTS-VAR successfully distinguishes references to different objects and generates accurate outputs based on prompts. For example, in the right second line, it eliminates conceptual mixtures and object disappearances. As illustrated in lines 3-10 on the right side, TTS-VAR also excels at determining counting numbers, positional relationships, and color attributes. Notably, in the last right line, our method generates a counterintuitive "green carrot," demonstrating its ability to separate objects from their natural attributes.
@@ Appendix H More Visualization Results
Appendix I Discussion on the Efficiency of Investing in Different Architectures
The efficiency of scaling across different model architectures remains an open question. In this section, we share our perspective on the potential benefits and limitations of scaling alternative architectures, such as VAR, in comparison to diffusion models.
First, different architectures possess distinct characteristics and advantages. Diffusion models, with their progressive denoising process and ability to model continuous latent spaces, generally achieve superior generative quality. By contrast, the VAR architecture more closely resembles the structure of large language models, which may be more conducive to unifying generation and understanding tasks [62]. Thus, depending on the specific objective, scaling efforts targeted at a given architecture may yield complementary benefits.
Second, we believe that current scaling strategies for diffusion models are still at an early stage. In contrast to the significant improvements demonstrated by our approach over Best-of-N (BoN), the search strategy introduced in [22] provides only marginal improvements over Random Search. We hypothesize that the multi-step denoising process inherent to diffusion models substantially enlarges the search space, thereby necessitating more sophisticated methodological designs. As a result, it remains difficult, based on current evidence, to make a definitive judgment on which scaling strategy or architectural choice is superior.
Appendix J Societal Impact
@@ Appendix I Discussion on the Efficiency of Investing in Different Architectures
Appendix K Limitation and Future Work
Though TTS-VAR shows significant improvement over the baseline and sets a new record, it still has two main limitations. First, TTS-VAR does not completely address the misalignment between text prompts and generated images. As indicated by the scores in Table˜7, there are still some failure cases, particularly in the Position item. Second, while TTS-VAR is based on a general coarse-to-fine process, its potential application to other coarse-to-fine models, such as autoregressive models that use 1-D tokenizers, remains unexplored. In the future, we will investigate the generation process more thoroughly, examining the reasons for failure and designing solutions to unlock further scaling potential. Additionally, we plan to assess the compatibility of TTS-VAR with other autoregressive coarse-to-fine models, including those utilizing 1-D tokenizers and hybrid architectures that combine diffusion models. These efforts aim to create a more robust scaling framework for text-to-image synthesis while enhancing the methodological transferability of coarse-to-fine paradigms.